OpenStudy (broskishelleh):

Task 3—Graphing Rational Functions Create a rational function with a linear binomial in both the numerator and denominator. Part 1. Graph your function using technology. Include the horizontal and vertical asymptotes and the x- and y-intercepts on your graph. Label the asymptotes and intercepts. Part 2. Show all work to identify the vertical asymptote, the x-intercepts, and the y-intercept. @Preetha @ganeshie8 @Abhisar @dan815 @e.mccormick @cram @paki @SolomonZelman @hartnn

OpenStudy (broskishelleh):

@Nurali @paki

OpenStudy (broskishelleh):

@shamil98

OpenStudy (broskishelleh):

@Compassionate

OpenStudy (broskishelleh):

@***[ISURU]***

OpenStudy (anonymous):

Well let's just say that f(x) = x+1 / x-1

OpenStudy (broskishelleh):

okay

OpenStudy (anonymous):

then you can graph that on your calculator vertical retricemtote is x= 1 since anything /0 is undefined.

OpenStudy (anonymous):

horizontal retricemtote is at y=1 because by calculus, limit of f(x) as x approaches +infinity and -inifinity is 1

OpenStudy (anonymous):

3. show work f(x) = (x+1)/ (x-1) f(1) = 1+1 / 1-1 f(1) = 2/0 = undefined. Therfore we have vertical retricemtote at x=1

OpenStudy (broskishelleh):

Can you please elaborate for me, and I am sorry if I disappear for a bit, my browser is broken

OpenStudy (broskishelleh):

Alright. got it. Thanks

OpenStudy (broskishelleh):

Is the answer f(x) = (x+1)/ (x-1) f(1) = 1+1 / 1-1 f(1) = 2/0 = undefined. Therfore we have vertical retricemtote at x=1

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

oh I mean horizontal retricemtote, you can look at that link :)

OpenStudy (anonymous):

My explantion, when you look for horizontal retricemtote, is the power of x equal on top and botton of fraction? x+1/x-1 seems like top and bottom have the same x^1

OpenStudy (anonymous):

Now you look at the coefficient infront of x y= x+ 1 / x-1 is same as 1x + 1 / 1x -1 Take in cosideration of x only when you find horizontal retricemtote of this graph y = 1x / 1x y= 1/1 y =1 So y = 1 is the horizontal retricemtote